You have just been offered a contract worth $ 1.12 million per year for 6 years.​ However, to take the​ contract, you will need to purchase some new equipment. Your discount rate for this project is 11.8 %. You are still negotiating the purchase price of the equipment. What is the most you can pay for the equipment and still have a positive NPV​?
Answer #1

Let Cash Flows in Year n be denoted by CFn

Amount paid for equipment = CF0

Given,

CF1 = $1.12 m
CF2 = $1.12 m
CF3 = $1.12 m
CF4 = $1.12 m
CF5 = $1.12 m
CF6 = $1.12 m

Discount Rate = r = 11.8%

Present Value = NPV = ΣCFn/(1+r)n

=> NPV = -CF0 + 1.12/(1+0.1) + 1.12/(1+0.1)2 + 1.12/(1+0.1)3 + 1.12/(1+0.1)4 + 1.12/(1+0.1)5 + 1.12/(1+0.1)6

=> NPV = -CF0 + 4.88

For NPV > 0,

-CF0 + 4.88 > 0

=> CF0 < 4.88 m

Hence, The maximum we can pay is $ 4.88 m